### v2

# Relevant or Random: Can LLMs Truly Perform Analogical Reasoning?

Abstract
Analogical reasoning is a unique ability of humans to address unfamiliar challenges by transferring strategies from relevant past experiences. One key finding in psychology is that compared with irrelevant past experiences, recalling relevant ones can help humans better handle new tasks Gentner and Smith (2012). Coincidentally, the NLP community has also recently found that self-generating relevant examples in the context can help large language models (LLMs) better solve a given problem than hand-crafted prompts Yasunaga et al. (2024). However, it is yet not clear whether relevance is the key factor eliciting such capability, i.e., can LLMs benefit more from self-generated relevant examples than irrelevant ones? In this work, we systematically explore whether LLMs can truly perform analogical reasoning on a diverse set of reasoning tasks. With extensive experiments and analysis, we show that self-generated random examples can surprisingly achieve comparable or even better performance, e.g., 4 performance boost on GSM8K with random biological examples. We find that the accuracy of self-generated examples is the key factor and subsequently design two improved methods with significantly reduced inference costs. Overall, we aim to advance a deeper understanding of LLM analogical reasoning and hope this work stimulates further research in the design of self-generated contexts.
Relevant or Random: Can LLMs Truly Perform Analogical Reasoning?
Chengwei Qin††thanks: Equal contribution, order decided by coin flip.†, Wenhan Xia11footnotemark: 1♣, Tan Wang11footnotemark: 1†, Fangkai Jiao†, Yuchen Hu†, Bosheng Ding†, Ruirui Chen, Shafiq Joty†♠ †Nanyang Technological University ♣Princeton University ♠Salesforce Research Institute of High Performance Computing (IHPC), Agency for Science, Technology and Research (A*STAR) 1 Fusionopolis Way, #16-16 Connexis, Singapore 138632, Republic of Singapore
1 Introduction
A hallmark of human intelligence is that they can solve novel problems by drawing analogy from relevant past experiences, a concept known as analogical reasoning in cognitive science (Vosniadou and Ortony, 1989). As indicated by the name, recalling previously acquired relevant experiences can facilitate humans to better tackle new tasks, whereas irrelevant ones are rarely beneficial and can even be distracting Gentner and Smith (2012). For instance, when faced with a novel math problem about determinants (e.g., calculating the value of a given fourth-order determinant), humans can resolve this by reflecting upon the methodology employed to ascertain the value of a third-order determinant, whereas biological knowledge (e.g., how the human body regulates its temperature) can generally be considered irrelevant.
With the recent advancements in scaling up model size and data, LLMs have demonstrated impressive zero-shot and few-shot performance across various reasoning tasks, especially, through advanced prompting methods like chain-of-thought (CoT) (Wei et al., 2022). Compared to common approaches such as zero or few-shot CoT (Zhou et al., 2022; Kojima et al., 2022; Wang et al., 2022b), Yasunaga et al. (2024) introduce LLM analogical reasoning, i.e., LLMs self-generate examples relevant to the query as context to better solve new problems; see Fig. 1 for an example. However, it remains unclear whether relevance is the key to eliciting such capability in LLMs. While several studies explore the influence of the relevance of demonstrations in in-context learning and CoT (Liu et al., 2022; Kim et al., 2022; Lyu et al., 2023; Chen et al., 2023; Yang et al., 2023; Wang et al., 2023; Yasunaga et al., 2024), none of them investigate whether self-generated relevant examples consistently outperform irrelevant ones in LLM analogical reasoning.
In this paper, to systematically assess the capability of LLMs to perform analogical reasoning, we conduct a series of ablation experiments on a variety of reasoning tasks including problems from GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and BIG-Bench Hard (BBH) (Suzgun et al., 2022). With extensive experiments, we aim to address the following two research questions:
Q1. Are self-generated relevant examples more beneficial to LLMs than random ones?
Q2. If not, what is the pivotal factor for LLMs’ performance in analogical reasoning?
To answer these questions, we empirically analyze the analogical reasoning abilities of GPT-3.5 (turbo) and Llama series models (Touvron et al., 2023). Surprisingly, experimental results show that prompting LLMs to self-generate random examples can achieve comparable or even better performance on certain tasks which is not in line with the key claim of analogical reasoning in Gentner and Smith (2012), indicating that LLMs cannot always perform analogical reasoning. As for Q2, we point out through controlled experiments that the key factor is the accuracy of self-generated examples. Informed by these findings, we design two approaches that can outperform existing methods with significantly reduced inference costs. Specifically, we ask LLMs to randomly generate a few problems and manually verify their correctness, then use this fixed set of problems as in-context learning demonstrations for all test samples. Consistent observations across different model types consolidate the conclusions. We summarize the major contributions of our work below:
To the best of our knowledge, we, for the first time, extensively assess the ability of LLMs to perform analogical reasoning and explore their counterintuitive behavior on certain tasks.
With extensive experiments and analysis, we demonstrate the effectiveness and limitations of different types of self-generated contexts.
2 Related Work
This work mainly explores whether LLMs can truly perform analogical reasoning. In light of this, we review two lines of research that form the basis of this work: chain-of-thought prompting and LLM analogical reasoning.
2.1 Chain-of-Thought Prompting
Chain-of-thought (CoT) prompting induces LLMs to generate intermediate reasoning steps before generating the final answer (Wei et al., 2022), greatly improving the reasoning capabilities of LLMs. Typical CoT prompting approaches include few-shot CoT (Wei et al., 2022; Zhou et al., 2022; Wang et al., 2022b; Li et al., 2022; Wang et al., 2022a), taking several labeled demonstrations of the reasoning process, and zero-shot CoT, comprising only instructions like “Let’s think step by step” (Kojima et al., 2022; Zelikman et al., 2022; Zhang et al., 2023). Other ongoing research on CoT has also explored (i) optimizing the demonstration selection (Fu et al., 2022; Lu et al., 2022; Qin et al., 2023), (ii) optimizing the quality of reasoning chains (Khot et al., 2022; Chen et al., 2022; Zhao et al., 2023; Shinn et al., 2023), and (iii) CoT in smaller language models (Magister et al., 2022; Ho et al., 2022; Fu et al., 2023).
2.2 LLM Analogical Reasoning
While few-shot CoT can provide more detailed reasoning guidance, it requires labeled examples which can be unavailable for a new task. To tackle this problem, Yasunaga et al. (2024) propose analogical prompting to guide LLMs to self-generate relevant exemplars as few-shot demonstrations, which is similar to analogical reasoning, i.e., humans can address new problems by drawing analogy from relevant past experience (Vosniadou and Ortony, 1989; Holyoak, 2012). In this work, we step forward to explore the intrinsic principle of LLM analogical reasoning. Specifically, we aim to investigate whether LLMs can authentically exhibit such reasoning capabilities and determine the extent to which the relevance of self-generated examples contributes to enhancing this process.
3 Methodology
We mainly follow the analogical prompting approach outlined in Yasunaga et al. (2024). Specifically, for a given target problem , analogical prompting introduces instructions like:
# Problem: []
# Relevant problems: Recall five relevant and diverse problems. For each problem, describe it and explain the solution.
# Solve the initial problem:
The goal is to induce LLMs to self-generate relevant examples, aiding them to solve the target problem via in-context learning. To ensure better performance and efficiency, several key technical decisions are made in Yasunaga et al. (2024):
The self-generated examples should be relevant and diverse, achieved through a specially designed instruction.
Generate relevant problems and the solution to the initial problem in one pass.
to self-generated examples perform the best.
In this work, we make minimal changes to the original prompt to guide LLMs to generate different types of irrelevant examples as context; see Fig. 2 for example prompts:
N/A: generate problems that are N/A (not applicable) to the initial problem.
: randomly generate examples of the same problem type (e.g., math).
: randomly generate examples of different problem types (e.g., any type except math).
: randomly generate biological problems.
Yasunaga et al. (2024) demonstrate that self-generating relevant examples can consistently outperform zero-shot CoT and few-shot CoT (hand-crafted examples or retrieved top- most similar training samples) on different tasks. Therefore, we do not include these two methods in our work. Interested readers can refer to the corresponding results and analysis in Yasunaga et al. (2024). In addition, we show prompts for different methods on all datasets in Section A.1.
4 Experiment
| Method | Temporal sequences | Logical deduction five objects | Reasoning about colored objects | Formal fallacies | Word sorting | Average |
| Relevant | 60.0 | 51.2 | 76.7 | 51.2 | 76.9 | 63.2 |
| N/A | 57.5 | 45.3 | 75.5 | 53.3 | 77.7 | 61.9 |
| 53.1 | 48.8 | 73.5 | 52.4 | 74.1 | 60.4 | |
| 44.3 | 44.8 | 72.4 | 51.2 | 69.2 | 56.4 | |
| 57.1 | 49.5 | 76.1 | 50.8 | 74.9 | 61.7 |
| Method | Task | ||
| GSM8K | MATH | Average | |
| Relevant | 71.5 | 33.3 | 52.4 |
| N/A | 75.5 | 36.1 | 55.8 |
| 75.1 | 36.3 | 55.7 | |
| 76.3 | 34.1 | 55.2 | |
| 75.3 | 34.6 | 54.9 | |
4.1 Experimental Setup
We construct the evaluation suite based on diverse reasoning-intensive tasks, including mathematical reasoning and other reasoning (e.g., logical and temporal reasoning) tasks:
Mathematical reasoning. We work with two commonly used datasets, GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021). For each dataset, we randomly sample 500 examples from the original test set and run experiments three times with different random seeds (resulting in different test samples).
Other reasoning. Following Yasunaga et al. (2024), we evaluate five reasoning tasks in BIG-Bench Hard (BBH) (Suzgun et al., 2022): temporal sequences (temporal reasoning), logical deduction five objects and reasoning about colored objects (logical reasoning), formal fallacies (deductive reasoning) and word sorting (symbolic reasoning). For each task, we use all test samples for evaluation and run experiments three times with different random seeds.
We mainly use GPT-3.5 (gpt-3.5-turbo) as the LLM and obtain all outputs from it with the temperature set to 0. We ask the LLM to self-generate 5 examples for GSM8K, 3 examples for MATH and BBH following Yasunaga et al. (2024).
4.2 Main Results
We now address the research questions asked in Section 1 with empirical results.
[…]
The results averaged over all random seeds are reported in Table 1 and Table 2; more detailed results for every seed are shown in Section A.2.
Self-generated relevant examples achieve the best average performance on BBH.
From the results in Table 1, we can observe that the superiority of self-generated relevant examples is empirically substantiated on BBH. Specifically, using relevant examples, denoted by ‘relevant’, outperforms other approaches on temporal and logical reasoning tasks. While it performs worse than ‘N/A’ on deductive and symbolical reasoning, it can still improve the accuracy by 1.3 on average compared to ‘N/A’.
However, the results on mathematical reasoning tasks are quite counterintuitive as described below:
Relevant examples do not guarantee better performance.
Different from BBH, all types of self-generated irrelevant examples consistently outperform relevant ones on both mathematical reasoning datasets, showing that LLMs cannot yet perform analogical reasoning on these tasks. Interestingly, when we use randomly generated biological examples (e.g., how the process of photosynthesis occurs in plants), they can yield about 2.5 better results on average compared to generating relevant math problems. Besides, ‘N/A’ achieves the best average result as it is the second-best on both datasets.
| Method | Precalculus | Intermediate Algebra | Algebra | Prealgebra | Counting & Probability | Geometry | Number Theory |
| Relevant | 10.4 | 9.8 | 51.8 | 56.8 | 22.1 | 24.2 | 37.0 |
| N/A | 9.1 | 15.7 | 55.5 | 61.0 | 28.7 | 25.8 | 34.2 |
| 12.3 | 17.6 | 54.4 | 60.6 | 25.4 | 25.8 | 34.9 | |
| 13.0 | 14.1 | 52.7 | 56.8 | 26.2 | 24.2 | 33.6 | |
| 13.0 | 12.2 | 53.0 | 59.2 | 28.7 | 25.8 | 32.2 |
Problems in MATH span various subjects and difficulty levels. To investigate whether the inferior performance of relevant examples on MATH is accidentally caused by certain categories, we further report the accuracy across different subjects and difficulty levels in Table 3 and Fig. 3. The consistent performance gap between ‘relevant’ and other methods across different problem categories demonstrates the inherent flaws of relevant examples, indicating that mathematical reasoning tasks exhibit different analogical reasoning paradigms from other reasoning tasks.
It might present challenges to prompt LLMs to accurately generate specific types of demonstrations. Therefore, given the unexpected results on mathematical reasoning tasks, one may wonder:
[…]
| Method | GSM8K | MATH | Average |
| Relevant | 0.54 | 0.41 | 0.48 |
| N/A | 0.19 | 0.28 | 0.24 |
| 0.30 | 0.20 | 0.25 | |
| 0.15 | 0.10 | 0.13 | |
| 0.06 | 0.11 | 0.09 | |
| Oracle | 0.65 | 0.63 | 0.64 |
| Relevant | N/A | ||
| Accuracy | 62.0 | 72.0 | 86.0 |
| Query: For how many ordered pairs where and are positive integers is | |
| Relevant | In a certain base, the sum of two three-digit numbers is . If the digits of one of the numbers are reversed, the sum becomes . What is the base of this number system? |
| N/A | What is the value of in the equation ? |
| In a bag, there are 5 red marbles, 3 blue marbles, and 2 green marbles. If you randomly pick 2 marbles from the bag without replacement, what is the probability that both marbles are red? | |
| How do you bake chocolate chip cookies? | |
| How does the process of photosynthesis occur in plants? | |
| Oracle | Find the number of ordered pairs of complex numbers such that . |
| Variant | GSM8K | MATH | ||||
| Relevant | N/A | Relevant | N/A | |||
| ICL | 71.2 | 73.8 | 72.0 | 37.0 | 39.8 | 39.2 |
| GPT4-Calibration | 75.2 | 75.6 | 75.6 | 44.4 | 41.2 | 40.0 |
| Random | 70.0 | 72.0 | 68.4 | 36.0 | 38.0 | 37.8 |
To quantitatively measure the relevance between the generated examples and the query, we compute the average cosine similarity between them. Following Zhang et al. (2023), we use the sentence transformer (Reimers and Gurevych, 2019) to encode all samples. For each method, the reported result is averaged across three seeds.
As observed from Table 4, relevant examples are much more semantically similar to the query than irrelevant ones and the relevance score of ‘relevant’ is more biased towards ‘oracle’ rather than ‘random’ or ‘N/A’, demonstrating that LLMs indeed follow instructions to generate specific types of demonstrations. Furthermore, we calculate the average similarity score between self-generated relevant examples and queries for BBH (0.46), which is slightly lower than the score of mathematical reasoning tasks (0.48). This result demonstrates that the difference in analogical reasoning performance between BBH (Table 1) and mathematical reasoning (Table 2) is not because LLMs can generate more relevant examples for BBH.
We provide a case study in Table 6 to delve deeper into the demonstrations of different methods. As we can notice, the example generated by ‘relevant’ is more related to the query as they both involve the mathematical concept ‘number bases’. In contrast, examples such as ‘What is the value of in the equation ?’ (N/A) or ‘How do you bake chocolate chip cookies?’ () are less relevant to the query. This comparison highlights once again that relevance may not be the key factor for analogical reasoning performance on mathematical reasoning tasks. To understand better the underlying reasons for the counterintuitive results, we then ask the following question:
[…]
Looking back at Table 6, an interesting observation is that the self-generated relevant example appears to be more difficult to solve than the irrelevant ones, regardless of whether they are math problems or not. Consequently, the accuracy of relevant examples may be lower. To verify this, we conduct a pilot experiment on MATH. Specifically, we randomly select 50 samples for different types of generated math problems, i.e., Relevant, N/A and , and manually evaluate their accuracy. We exclude other methods as it is difficult to define the ‘accuracy’ of the examples they generate. From the results in Table 5, we can observe that while the examples generated by ‘relevant’ are more related to the test query, they are less accurate, raising the question whether the performance of different approaches on mathematical reasoning tasks is strongly correlated with the accuracy of self-generated examples.
Proxy Approaches
However, as the accuracy of the examples located at the output cannot be directly controlled, we meticulously design a variant called ICL, which extracts the generated examples from the model output as in-context learning (ICL) demonstrations and combines them with the query as input to LLMs, as a proxy for the original method. We also consider the following two variants: (a) GPT4-Calibration which replaces the answers of demonstrations in ICL with GPT4-generated answers, and (b) Random changes the answers of demonstrations in ICL to random numbers. Our manual verification confirmed that GPT4-generated answers were mostly accurate. We conduct this experiment on GSM8K and MATH with GPT-3.5 as the LLM reasoner.
| Method | Task | ||
| GSM8K | MATH | Average | |
| Relevant | 71.5 | 33.3 | 52.4 |
| N/A | 75.5 | 36.1 | 55.8 |
| 75.1 | 36.3 | 55.7 | |
| 76.3 | 34.1 | 55.2 | |
| 75.3 | 34.6 | 54.9 | |
| 75.7 | 36.8 | 56.3 | |
| 77.9 | 34.9 | 56.4 | |
From the results of different variants reported in Table 7, we can see that increasing the accuracy of generated examples can indeed improve the performance: GPT4-Calibration consistently outperforms ICL by incorporating more accurate answers. In contrast, random always performs the worst among all variants. Therefore, the key factor influencing the performance on mathematical reasoning is the accuracy of self-generated examples rather than their relevance.
It is worthwhile to note that while several papers explore how the correctness of demonstration answers influences in-context learning (Min et al., 2022; Yoo et al., 2022; Wei et al., 2023; Pan et al., 2023; Kossen et al., 2024), our work differs from them in the following aspects: (i) The examples in our work are generated by LLMs rather than real data from NLP benchmarks, i.e., randomly sampled from the training set. In addition, there are rationales (CoT) in self-generated examples, which are different from the input-label format of in-context learning investigated in these papers; and (ii) These studies mainly evaluate in-context learning on different classification or multi-choice datasets, i.e., the output space is a finite set. In contrast, we are evaluating mathematical reasoning tasks, where the output space is infinite.
Given the above findings, a natural question is:
[…]
We refer to these two methods as and , and conduct experiments with them on GSM8K and MATH (see Fig. 4 for example prompts and outputs for generating math problems). Detailed prompts and outputs for different methods are provided in Section A.3. Following the original setting, we ask the LLM to randomly generate 5 examples for GSM8K and 3 examples for MATH. As observed from Table 8, ensuring the accuracy of self-generated examples does lead to better performance regardless of the problem type. and achieve similar average performance, once again demonstrating that relevance does not matter (see Section A.4 for more analysis on relevance). Moreover, both ICL variants only need to generate examples once, which significantly reduces the inference cost and further demonstrates their superiority.
| Method | Task | ||
| GSM8K | MATH | Average | |
| 75.7 | 36.8 | 56.3 | |
| 73.8 | 36.2 | 55.0 | |
| Relevant | N/A | |||
| 71.2 | 74.9 | 75.3 | 75.9 | 74.3 |
| Method | Relevant | N/A | |||||
| Llama-2-70b-Chat | 45.1 | 51.4 | 50.9 | 54.3 | 47.1 | 55.5 | 56.1 |
| Llama-3-8B-Instruct | 69.5 | 72.3 | 72.6 | 74.1 | 73.5 | 75.8 | 76.8 |
| Variant | Method | ||
| Relevant | N/A | ||
| ICL | 56.2 | 58.2 | 58.6 |
| GPT4-Calibration | 60.8 | 61.0 | 60.8 |
| Random | 53.2 | 54.0 | 59.6 |
4.3 Further Analysis
Repeating Problems
While generating a few accurate problems as ICL demonstrations can achieve better performance, a bolder idea might be to generate one problem and repeat it multiple times as few-shot demonstrations for ICL. To investigate this, we randomly select a generated math problem and repeat it to perform ICL, denoted by . From the results shown in Table 9, we can see that consistently performs worse than on both datasets, indicating that the diversity of generated problems also matters.
Robustness to Prompt Formats
To verify the robustness of different methods to prompt formats, we experiment with a new prompt paraphrased from the original one by GPT-4 and present the results on GSM8K in Table 10. We also observe better performance with irrelevant examples than relevant ones, showing the robustness.
Generalization to Open-source LLMs
Our experiments and analysis so far used GPT-3.5 as the LLM, which is closed-source and gets updated over time. To verify whether the observations and conclusions are consistent across different models and additionally for reproducibility, we extend the experiments to Llama-2-Chat (Touvron et al., 2023). Specifically, we use vLLM (Kwon et al., 2023) to serve a Llama-2-70b-Chat model for experiments and report the results of different methods/variants on GSM8K in Table 11 and Table 12. We can draw similar observations: (i) self-generated relevant examples underperform all types of irrelevant ones, (ii) ‘GPT4-Calibration’ consistently outperforms the other two variants, and (iii) and perform better than other approaches, demonstrating that the conclusions can be generalized to different models.
We further conduct experiments with the latest Llama-3-8B-Instruct model. The results reported in Table 11 demonstrate the generalizability of the conclusions across different model scales. In addition, since investigating analogical reasoning requires LLMs to self-generate different types of problems, we only experiment with instruction-tuned LLMs to ensure that they can follow the given instructions.
Different Numbers of Demonstrations
While we mainly follow the setting in Yasunaga et al. (2024) to ask the LLM to generate examples for GSM8K, we further investigate the effect of the number of demonstrations. Specifically, we conduct controlled experiments with and report the results in Table 13. We can observe that irrelevant examples consistently outperform relevant ones across different numbers of demonstrations, emphasizing their effectiveness.
| Number | Relevant | N/A | |||
| 3 | 73.1 | 77.3 | 75.0 | 75.3 | 75.5 |
| 5 | 71.5 | 75.5 | 75.1 | 76.3 | 75.3 |
5 Conclusion
In this work, we have systematically assessed the capability of LLMs to perform analogical reasoning. We have identified key research questions and empirically analyzed a representative set of LLMs on a diverse collection of reasoning tasks. Extensive experimental results and analysis show that LLMs cannot always perform analogical reasoning and the key influencing factor is the accuracy of self-generated examples rather than their relevance. Given these findings, we have designed two ICL-based approaches with better performance and significantly reduced inference costs. In the future, we would like to investigate additional analogical prompting methods to generate more accurate examples.
Limitations
This work has several limitations. First, due to the inference cost of ChatGPT111experiments done between 12/2023 and 05/2024, we conduct experiments on subsets of the test data for mathematical reasoning tasks. Besides, we include 3 datasets requiring different reasoning capabilities in this work. A further improvement could be to explore more diverse types of tasks.
Appendix A Appendix
A.1 Prompts for Different Methods
A.2 Detailed Results for Different Random Seeds
| Seed | GSM8K | MATH | ||||||||
| Relevant | N/A | Relevant | N/A | |||||||
| 42 | 71.8 | 76.6 | 73.2 | 74.0 | 74.0 | 37.4 | 42.2 | 41.6 | 39.0 | 39.2 |
| 100 | 71.2 | 75.2 | 75.2 | 75.8 | 74.8 | 29.0 | 30.6 | 32.6 | 29.4 | 31.2 |
| 1000 | 71.4 | 74.8 | 77.0 | 79.2 | 77.0 | 33.6 | 35.6 | 34.6 | 34.0 | 33.4 |
| Average | 71.5 | 75.5 | 75.1 | 76.3 | 75.3 | 33.3 | 36.1 | 36.3 | 34.1 | 34.6 |
| Seed | Temporal sequences | Logical deduction five objects | Reasoning about colored objects | Formal fallacies | Word sorting | Average | |
| 42 | Relevant | 58.0 | 52.8 | 76.0 | 50.4 | 77.2 | 62.9 |
| N/A | 56.4 | 44.8 | 77.6 | 54.0 | 76.8 | 61.9 | |
| 52.4 | 48.8 | 74.8 | 51.6 | 72.8 | 60.1 | ||
| 43.2 | 46.8 | 74.0 | 52.4 | 67.6 | 56.8 | ||
| 56.8 | 52.0 | 74.0 | 52.0 | 76.4 | 62.2 | ||
| 100 | Relevant | 58.4 | 50.8 | 78.4 | 51.2 | 76.8 | 63.1 |
| N/A | 55.2 | 46.0 | 74.8 | 52.8 | 79.2 | 61.6 | |
| 50.8 | 48.4 | 73.6 | 53.2 | 75.2 | 60.2 | ||
| 46.4 | 46.8 | 72.8 | 50.0 | 70.4 | 57.3 | ||
| 58.0 | 48.4 | 78.4 | 51.2 | 73.6 | 61.9 | ||
| 1000 | Relevant | 63.6 | 50.0 | 75.6 | 52.0 | 76.8 | 63.6 |
| N/A | 60.8 | 45.2 | 74.0 | 53.2 | 77.2 | 62.1 | |
| 56.0 | 49.2 | 72.0 | 52.4 | 74.4 | 60.8 | ||
| 43.2 | 40.8 | 70.4 | 51.2 | 69.6 | 55.0 | ||
| 56.4 | 48.0 | 76.0 | 49.2 | 74.8 | 60.9 | ||
A.3 Prompts and Outputs for Example Generation
A.4 Guided Problem Generation
In addition to random problem generation in Section 4.2-Q2-1, we further investigate guided problem generation. Specifically, we randomly select 5 training samples to guide LLMs to self-generate relevant math problems. We then manually verify their correctness and use this fixed set of problems as ICL demonstrations for experiments. The performance of this approach (56.1) is slightly lower than that of (56.3), verifying that relevance is not the key influencing factor.